Hi ....i have a question from permution and combination.


What is the value of 1*1! + 2*2! + 3!*3! + ............ n*n!,

where n! means n factorial or n(n-1)(n-2)...1


plz reply !!!!!


 






Grade:

1 Answers

raj kaushik
15 years ago

1*1! = (2 -1)*1! = 2*1! - 1*1! = 2! - 1!

2*2! = (3 - 1)*2! = 3*2! - 2! = 3! - 2!

3*3! = (4 - 1)*3! = 4*3! - 3! = 4! - 3!
..
..
..

n*n! = (n+1 - 1)*n! = (n+1)(n!) - n! = (n+1)! - n!

Summing up all these terms, we get (n+1)! - 1!